# Ginkeepaspectratio=true,clip=true,draft=false

### v1

@@ Abstract
We present LionsOS, an operating system for security- and safety-critical embedded systems. LionsOS is based on the formally verified seL4 microkernel and designed with verification in mind. It uses a static architecture and features a highly modular design driven by strict separation of concerns and a focus on simplicity. We demonstrate that LionsOS outperforms Linux.
@@ 1 Introduction
Safety- and security-critical systems, such as aircraft, autonomous cars, industrial control systems or defence systems, require a highly dependable operating system (OS). The complexity (and code size) of these embedded/cyber-physical systems keeps growing, and it is unavoidable to have highly critical (and, presumably, highly assured) functionality co-exist with less critical (and less trustworthy) functionality. These systems are therefore mixed criticality systems (MCS), where the correct operation of a critical component must not depend on the correctness of a less critical component [Barhorst et al., 2009; Burns and Davis, 2017].
Unlike more traditional MCS, which are mostly concerned with (spatial and temporal) integrity and availability, the mixed-criticality requirement must also extend to confidentiality: Assume an untrusted component of the system is compromised by an attacker, and this component can be used to obtain the system’s encryption key used to validate over-the-air software upgrades, then the attacker could replace critical components and thus completely compromise the system.
@@
The core requirement such critical systems impose on the OS is strong temporal and spatial isolation. As such, the seL4 microkernel [seL4 Foundation, 2021a] seems a perfect foundation: seL4 has undergone extensive formal verification, including proofs of confidentiality and integrity enforcement, proof of implementation correctness and proofs that the binary has the same semantics as the verified C code, taking the compiler out of the trusted computing base (TCB) [Klein et al., 2014]. There is furthermore a MCS version of seL4 which provides the temporal isolation properties required by real-time systems [Lyons et al., 2018], verification of the MCS variant is currently in progress [seL4 Foundation, 2024].
@@
These features have resulted in some deployment success, including autonomous military aircraft [Cofer et al., 2018] and, more recently, commercial electric cars [Qu, 2024]. Yet more than a decade after its verification was completed, seL4 is not as widely deployed as might have been expected.
The core reason behind this slow uptake is the low-level nature of seL4, which goes beyond the already low-level nature of other microkernel APIs. For example, seL4 makes all memory management (including for the kernel) a user-level responsibility [Elkaduwe et al., 2008]; while this is a core enabler of reasoning about isolation, the developer experiences it as a foot gun. Together with a lack of support for writing device drivers, this makes seL4 the “assembly language of operating systems”. The upshot is that it requires an unusual level of expertise to build functioning and performant systems on seL4 – the kernel is largely unapproachable by industrial developers of critical systems.
@@
In short, to benefit from seL4’s provable isolation enforcement, developers of critical systems need an actual operating system, providing appropriate abstractions, such as processes, files and network connections, while retaining (and extending) the isolation guarantees provided by seL4. We propose to address this need with LionsOS, designed to meet the following aims:
Fast:
LionsOS performance should be competitive with that of mainstream systems, such as Linux;
Secure:
LionsOS should, by design, not only leverage seL4’s verified isolation properties to protect applications from each other, but make an eventual formal verification of LionsOS itself tractable;
Adaptable:
LionsOS should be deployable on most critical embedded, IoT and cyberphysical systems.
These aims are somewhat in conflict: adaptability and verifiability call for a highly modular design. While such a design is natural for a microkernel-based system, the cost of context switches could undermine performance [Chen et al., 2024]. However, in this paper we show that a highly principled (Section 3), modular design (Section 4), combined with an implementation focused on simplicity (Section 5) can result in a system that does meet the performance aim (Section 6).
@@ 2.1 Verification and scalability
The formal verification of seL4 demonstrated that it is possible to prove the implementation correctness of real-world systems of considerable complexity. However, the cost was high: about 12 person years of non-recurring engineering for 8,500 source lines of code (SLOC), and an estimated cost of \$350/SLOC [Klein et al., 2009]. While potentially justified for a stable, foundational piece of infrastructure, this cost is too high for most systems, including an OS that will be significantly larger than seL4.
The seL4 project delivered another key insight: Verification effort scales with the square of the specification size [Matichuk et al., 2015]. This implies that there could be a large scalability benefit from keeping things small and simple, i.e. a highly modular design, where each component has a narrow and simple specification, and where module boundaries are enforced by seL4, making it possible to verify modules independently of each other.
@@ 2.2 Modularity in operating systems
These systems were plagued by poor performance, resulting in moving functionality back into the kernel [Welch, 1991], which did not prevent expensive debacles, such as IBM’s ill-fated, Mach-based Workplace OS [Fleisch et al., 1998].
Almost all microkernel-based OSes exhibited course-grained modularity, typically at the level of major subsystems such as file service, networking and process management  [Rawson, 1997; Whitaker et al., 2002; Härtig et al., 2005; Herder et al., 2006; Qubes, ], too large for verification. Nevertheless, even the most recent work argues that the cost of crossing module boundaries is too high, resulting in co-locating services into even more course-grained isolation domains [Chen et al., 2024].
The Flux OSKit [Ford et al., 1997] was an early design featuring a more fine-granular design. Performance comparison to Linux and FreeBSD showed a 13% degradation in networking throughput and a 45% increase in latency. SawMill was an ambitious project aiming to break up Linux into components isolated by a microkernel, file-system benchmarks showed a throughput degradation of about 18%. No CPU load values are reported for any of these systems, but the degradation in achieved throughput indicates a significant increase in per-packet processing cost.
Genode (formerly called Bastei) [Feske and Helmuth, 2007; Feske, 2015] features a modular design aiming for assurance and explicitly prioritising this over performance (we could not find any published performance data). Its implementation in C++ will prevent a complete formal verification for the foreseeable future.
An alternative is modularisation enforced by programming languages (as opposed to address-space isolation mediated by a microkernel), as pioneered by SPIN [Bershad et al., 1995], and later adopted by Singularity [Fähndrich et al., 2006] and RedLeaf [Narayanan et al., 2020]. These systems generally exhibit lower performance than mainstream OSes. Furthermore, as their security relies on type-safety enforcement by the programming language, they require the whole OS to be implemented in that language. Inevitably this requires unsafe escapes for dealing with hardware. More importantly, this rules out re-using code from mainstream OSes.
@@ 3 LionsOS Design Principles
Given this experience, what makes us think we can meet all our aims from Section 1?
We observe that a commonality of these earlier systems is a significant complexity in design and implementation. We posit that the key to meeting our aims is a strict application of the time-honoured KISS Principle [Wikipedia, 2001]. Following this high-level principle, we aim for a highly modular design which incorporates the following secondary principles:
@@
Each module has one and only one purpose (as far as feasible). Furthermore, a particular concern (e.g. the traffic-shaping policy of a network subsystem) should be fully contained in a single module.
@@
Each module only has the access rights it needs, not more. While not strictly a consequence of KISS, this time-honoured security principle of Saltzer and Schroeder [1975] will simplify reasoning about security and safety of the system.
@@
Given the scalability and complexity limits of verification (Section 2.1), this principle calls for keeping module interfaces narrow and module implementations simple.
@@
This is arguably the most controversial principle, as it calls for tailoring each (resource) policy to the system’s particular use case, in order to simplify the policy implementation.
@@
The common thread of the embedded domain is that it can be served with a static system architecture, i.e. a set of components that is known at configuration time. Note that this does not mean that the system is fully static, it can still support late loading of components, dynamic component updates, and place-holders for components that can be instantiated with programs not known at build time. Dependable embedded systems generally cannot over-commit resources, which is what makes the static architecture work. We are yet to see a realistic use case in the embedded space that cannot be addressed with a static architecture.
@@
But the embedded system’s defined set of resources implies that, at least for the critical systems we are targeting, the designer also has (or should have) a clear idea of how they should be managed. The more tailored the policy is to the use case, the simpler its implementation, and the easier it is to assure (formally or informally) that it matches requirements.
This principle of use-case specific policies is arguably the clearest departure from conventional approaches. OSes tend to be designed to adapt to changing application scenarios with no or minimal code changes. This naturally leads to a (whether conscious or not) desire to provide universal policies. Of course, no policy is truly universal, and sooner or later it will encounter a use case where the existing policy behaves pathologically, resulting in attempts to generalise it. This approach is a massive driver of complexity: For example, the Linux scheduler contains five scheduling classes, each of which has one or more per-thread tuning parameters; the scheduler has grown from around 11,kSLOC in 2011 (version 3.0) to over 30 kSLOC today (version 6.12). Furthermore, the approach frequently leads to optimising particular “hot” use cases at the expense of overall performance [Ren et al., 2019].
@@
An illustrative example is network traffic shaping: If multiple clients of a network interface overload the interface, there really are a small number of obvious policies to choose from: Clients may be given a priority, their bandwidth may be limited, or they may be served round-robin. And in an embedded system, it is usually obvious which one is appropriate, and this is unlikely to change for the lifetime of the system.
@@
An obvious concern is how the fine-granular modularity affects our performance aim, as this necessarily leads to high context-switching rates. Given there is a cost to every context switch (depending on the architecture, 400–600 cycles for seL4 [seL4 Foundation, 2021b]), this has the potential to make the system slow [Chen et al., 2024]. We will examine the performance impact in Section 6.
@@ 4 LionsOS Design
Like most OSes, LionsOS consists mostly of I/O subsystems (device drivers, protocol stacks, file systems) and resource management. The latter part is particularly small in LionsOS because of its static architecture, which reduces resource management to some simple policy modules that are not only use-case specific (as per our principle) but also local in their nature (e.g. shaping traffic of a network interface).
Some global resource management is required but not yet fully designed (nor implemented), such as core management (off- and on-lining processor cores based on overall CPU load). We will discuss this in Section 5.3.
@@ 4.1.1 Device drivers
Applying separation of concerns, we reduce the purpose of a device driver to translate between a hardware-specific device interface and a hardware-independent device class interface. Hence an Ethernet driver abstracts the specific NIC as a generic Ethernet device.
Unsurprisingly, the Ethernet device-class interface (i.e. the driver’s OS interface) looks similar to that of an actual Ethernet NIC, with some differences that help simplify its use. NICs typically use ring buffers in DMA memory to pass references to DMA buffers from and to the driver; each ring buffer entry contains a pointer to a buffer in the DMA region, together with some meta-data (indicating whether a buffer contains valid data). A NIC usually references two such ring buffers, one for transmit (Tx) and one for receive (Rx) data.
To simplify the OS interface, we separate queues for buffers containing valid data from those that do not. This means that the SW side of the driver has four queues:
@@
references buffers provided by the OS to be sent to the network;
@@
These queues are allocated in driver metadata regions, for Ethernet there is one for Tx data (containing the TxA and TxF queues) and a separate one for Rx data (RxA and RxF queues). These are “normal” memory invisible to the device (i.e. not accessed by DMA).
Note that the driver only handles pointers to DMA buffers, it has no need to access the actual data. We make this explicit by separating the data region, which contains the buffers to be filled/emptied by the device, from the device metadata region, which contains the ring buffers pointing to the data buffers. In addition there is the device control region, which maps the device registers for memory-mapped I/O. Data and device metadata regions are memory regions accessed by the device via DMA.
As the driver has no need to access the I/O data, it does not have the data region mapped into its address space, in line with the principle of least privilege. The device control region is mapped to the driver uncached.
The right half of Figure 1 shows the memory regions relevant to the driver. For the Ethernet device class, we treat the Tx and Rx paths separately, hence the driver’s OS interface has two metadata regions, one for transmit and one for receive.
@@ 4.1.2 Virtualisers
While the driver abstracts the device hardware, it does not deal with address translation, nor with sharing the device between multiple clients. This is the purpose of a separate virtualiser (Virt) component, which:
  • shares a physical device between multiple clients, presenting each with a virtual device;
@@
  • performs cache management (flushing/invalidating) where needed (the x86 architecture keeps caches coherent with DMA and thus does not require this functionality).
@@
The Virt replicates its driver interface at the client side, meaning its client-side interfaces look exactly like its device interface. Specifically it has a metadata region (Tx V MD, Rx V MD) that structurally replicates the respective device metadata region. The key difference is that while the device MD regions use I/O addresses for referring to data buffers, the Virt MD regions refers to data buffers by offsets from the beginning of the respective data region, thus making the Virt’s address-translation task independent of client virtual address-space layout.
The RxVirt needs to inspect the headers of incoming packets and thus requires the data region to be mapped (R/O). The TxVirt does not need to access the data region directly, but may need to have it mapped in its address space to perform cache management. For example, the Arm architecture performs cache operations on virtual address ranges, so the TxVirt needs the data region to be mapped into its address space. Arm does not have cache-coherent DMA, so memory that will be transferred to a device using DMA has to be cache-cleaned before DMA occurs. Likewise, the RxVirt on Arm needs to invalidate caches after DMA into its buffers. These mappings are indicated in the left half of Figure 1.
The TxVirt must implement a traffic shaping policy if its clients generate load that exceeds the NIC’s Tx capacity. In line with our principle of use-case specific policy, the Tx policy is as simple as the specific use-case allows. It is typically one of round-robin, priority-based or bandwidth limiting.
The RxVirt may only require a simple policy: what to do when data arrives for a client whose RxA queue is full, possible choices are to block or (more likely) discard the packet and return the buffer to the driver’s RxF queue. We generally avoid this this need by ensuring that all client queues are large enough to hold all available Rx buffers, starving the device of buffers if the clients fail to process input fast enough – this leads to the NIC dropping packets under overload without wasting CPU cycles, and leaves the RxVirt policy-free.

### v2

@@ Abstract
We present LionsOS, an operating system for security- and safety-critical embedded systems. LionsOS is based on the formally verified seL4 microkernel and designed with verification in mind. It uses a static architecture and features a highly modular design driven by strict separation of concerns and a focus on simplicity. We demonstrate that LionsOS achieves excellent performance on system-call intensive workloads.
@@ 1 Introduction
Safety- and security-critical systems, such as aircraft, autonomous cars, medical devices, industrial control or defence systems, require a highly dependable operating system (OS). The complexity (and code size) of these embedded/cyber-physical systems keeps growing, and it is unavoidable to have highly critical (and, presumably, highly assured) functionality co-exist with less critical (and less trustworthy) functionality. These systems are therefore mixed criticality systems (MCS), where the correct operation of a critical component must not depend on the correctness of a less critical component [Barhorst et al., 2009; Burns and Davis, 2017].
Traditional MCS are concerned with (spatial and temporal) integrity and availability. In modern systems the mixed-criticality requirement must also extend to confidentiality: Many cyberphysical systems, such as drones or medical devices, process sensitive data that must be protected.
@@
The core requirement such critical systems impose on the OS is therefore strong temporal and spatial isolation. As such, the seL4 microkernel [seL4 Foundation, 2021b] seems a perfect foundation: seL4 has undergone extensive formal verification, including proofs of confidentiality and integrity enforcement, proof of implementation correctness and proofs that the binary has the same semantics as the verified C code, taking the compiler out of the trusted computing base (TCB) [Klein et al., 2014]. It is the only formally-verified kernel that uses fine-grained access control through capabilities [Dennis and Van Horn, 1966]. There exists an MCS version of seL4 that provides the temporal isolation properties required by real-time systems [Lyons et al., 2018], verification of the MCS variant is currently in progress [seL4 Foundation, 2024].
@@
These features have resulted in some real-world deployments, including autonomous military aircraft [Cofer et al., 2018] and, more recently, commercial electric cars [Qu, 2024]. Yet more than a decade after seL4’s verification was completed, it is not widely deployed.
The core reason behind this slow uptake seems to be the low-level nature of seL4, even lower than for most other microkernels. For example, seL4 makes management of all of a system’s physical memory (including the kernel’s) a user-level responsibility [Elkaduwe et al., 2008]; while this is a core enabler of reasoning about isolation, managing memory requires data structures (e.g. page tables) which require memory – a foot gun for developers.
seL4 can be said to be the “assembly language of operating systems”. To build functional, performant systems on seL4 requires deep expertise – the kernel is largely unapproachable by industrial developers of critical systems, resulting in a number of seL4-based deployment projects apparently abandoned a few years after they started.
@@
In short, to benefit from seL4’s provable isolation enforcement, developers of critical systems need an actual operating system, providing appropriate abstractions, such as processes, files and network connections, while retaining (and extending) the isolation guarantees provided by seL4. We furthermore posit that such a system must be highly modular, in order to make best use of seL4-provided isolation for minimising the impact of faults, simplify identification and elimination of bugs, and enable end-to-end formal verification (eventually – for now we leave verification out of scope but aim for a verification-friendly design.)
Fine-grained modularity results in many context switches and IPC operations; this overhead has traditionally been considered the Achilles heel of microkernels [Bershad, 1992]. Even very recent microkernel-based work co-locates services in larger modules and migrates functionality back into the kernel in order to achieve performance competitive with the monolithic Linux system [Chen et al., 2024].
While such performance challenges may be real for a general-purpose OS system (eg. for smartphones), we aim to achieve a performant yet highly-modular OS for the embedded/cyberphysical space.
Our contribution is to demonstrate
  • LionsOS, the first highly-modular microkernel-based OS that achieves performance at par or better than traditional monolithic designs;
  • which is enabled by a simplicity-oriented design employing use-case-specific policies.
We present the principles of our approach (Section 3), discuss the resulting design (Section 4) and its implementation focussed on simplicity (Section 5). We evaluate LionsOS performance against Linux and microkernel-based OSes and find that on context-switch intensive loads, which can be expected to show the highest modularity-imposed overheads, LionsOS outperforms all systems we compare against (Section 6).
@@ 2.1 Verification and scalability
The formal verification of seL4 demonstrated that it is possible to prove the implementation correctness of real-world systems of considerable complexity. However, the cost was high: about 12 person years of non-recurring engineering for 8,500 source lines of code (SLOC), and an estimated cost of \$350/SLOC [Klein et al., 2009]. While potentially justified for a stable, foundational piece of infrastructure, this cost is too high for most systems, especially a complete OS that will be significantly larger than seL4.
The seL4 project delivered another key insight: Verification effort scales with the square of the specification size [Matichuk et al., 2015]. This implies that there could be a large scalability benefit from keeping things small and simple, i.e. a highly modular design, where each component has a narrow interfaces and simple specification, and where module boundaries are enforced by seL4, making it possible to verify modules independently of each other.
@@ 2.2 Modularity in operating systems
These systems were plagued by poor performance, and functionality moved back into the kernel [Welch, 1991]. This did not prevent expensive debacles, such as IBM’s ill-fated, Mach-based Workplace OS [Fleisch et al., 1998].
Almost all microkernel-based OSes exhibited course-grained modularity, typically at the level of major subsystems such as file service, networking and process management  [Rawson, 1997; Whitaker et al., 2002; Härtig et al., 2005; Herder et al., 2006; Qubes, 2010], making them too large for verification. Nevertheless, even the most recent work argues that the cost of crossing module boundaries is too high, resulting in a drastically-expanded kernel of 90 kSLOC (vs. seL4’s 10 kSLOC) and co-locating services into even more course-grained isolation domains [Chen et al., 2024].
The Flux OSKit [Ford et al., 1997] was an early design featuring a more fine-grained design. Performance comparison to Linux and FreeBSD showed a 13% degradation in networking throughput and a 45% increase in latency. SawMill [Gefflaut et al., 2000] was an ambitious project aiming to break up Linux into components isolated by a microkernel; file-system benchmarks showed a throughput degradation of about 18%. No CPU load values are reported for any of these systems, but the degradation in achieved throughput indicates a significant increase in per-packet processing cost.
Genode (formerly Bastei) [Feske and Helmuth, 2007; Feske, 2015] features a modular design, explicitly prioritising assurance over performance (we could not find any published performance data, although we evaluate against it in Section 6.3.1). Its implementation in C++ will prevent a complete formal verification for the foreseeable future. THINK [Fassino et al., 2002] is a component system for building kernels rather than an OS with isolated components. Systems like TinyOS [Levis et al., 2005], Tock [Levy et al., 2017] and Tinkertoy [Wang and Seltzer, 2022] are for microcontrollers without memory protection and as such unsuitable for MCS.
An alternative to enforcing modularity by address-space isolation is enforcing it through the programming language, as pioneered by SPIN [Bershad et al., 1995], and later adopted by Singularity [Fähndrich et al., 2006] and RedLeaf [Narayanan et al., 2020]. These systems generally exhibit lower performance than mainstream OSes. Furthermore, as their security relies on type-safety enforcement by the programming language, they require the whole OS to be implemented in that language. Inevitably this requires unsafe escapes for dealing with hardware. More importantly, this rules out re-using code from mainstream OSes.
@@
What is common to these earlier systems is their complexity, among others driven by the desire for code or API re-use leading to design compromises. Our core take-away is that to make modular OSes work, a clean, principled from-scratch approach is needed. Re-use, while desirable, should be subject to clean design principles, rather than compromising them.
@@ 3 LionsOS Design Principles
Given this experience, what makes us think we can meet our performance aim from Section 1 with a modular system?
We observe that a commonality of these earlier systems is a significant complexity in design and implementation. We posit that the key to meeting aim is a strict application of the time-honoured KISS Principle [Wikipedia, 2001]. Following this high-level principle, we aim for a highly modular design which incorporates the following secondary principles:
@@
Each module has one and only one purpose (as far as feasible). Furthermore, a particular concern (e.g. the network traffic-shaping policy) should be fully contained in a single module.
@@
Each module only has the access rights it needs, not more. While not a consequence of KISS, this time-honoured security principle of Saltzer and Schroeder [1975] simplifies reasoning about security and safety.
@@
Module interfaces are narrow and module implementations simple, to make verification scalable (Section 2.1).
@@
Likely the most controversial, this principle calls for tailoring each (resource) policy to the system’s particular use case, in order to simplify the policy implementation.
@@
The common thread of the embedded domain is that it can be served with a static system architecture, i.e. a set of components that is known at configuration time. This does not mean that the system is fully static – it can support late loading of components, dynamic component updates, and place-holders for components that are loaded with programs not known at build time. Dependable embedded systems generally cannot over-commit resources, which is what makes the static architecture work. We are yet to see a realistic use case in the embedded space that cannot be addressed with a static architecture. Note that the static system architecture does not prevent a subsystem, such as a virtual machine (VM) from managing a subset of resources dynamically.
@@
The embedded system’s defined set of resources implies that, at least for the critical systems we are targeting, the designer has (or should have) a clear idea of how they should be managed. The more tailored the policy is to the use case, the simpler its implementation, and the easier it is to assure (formally or informally) that it matches requirements.
This principle of use-case specific policies is arguably the clearest departure from conventional approaches. OSes tend to be designed to adapt to changing application scenarios with no or minimal code changes. This naturally leads to a (conscious or not) desire to provide universal policies. Of course, no policy is truly universal, and sooner or later it will encounter a use case where the existing policy behaves pathologically, resulting in attempts to generalise it. This approach is a massive driver of complexity: For example, the Linux scheduler contains five scheduling classes, each of which has one or more per-thread tuning parameters; the scheduler has grown from around 11,kSLOC in 2011 (version 3.0) to over 30 kSLOC today (version 6.12). Furthermore, the approach frequently leads to optimising particular “hot” use cases at the expense of overall performance [Ren et al., 2019].
@@
An illustrative example is network traffic shaping: If multiple clients of a network interface overload the interface, there are only a small number of obvious policies to choose from: Clients may be given a priority, their bandwidth may be limited, they may be served round-robin, or certain protocols might be prioritised. In an embedded system, it is usually obvious which one is appropriate, and this is unlikely to change for the lifetime of the system.
@@
An obvious concern is how the fine-granular modularity affects our performance aim, as modularity necessarily leads to high context-switching rates. Given there is a cost to every context switch (depending on the architecture, 400–600 cycles for seL4 [seL4 Foundation, 2021a]), this has the potential to make the system slow [Chen et al., 2024]. We will examine the performance impact in Section 6.
@@ 4 LionsOS Design
LionsOS is based on seL4, and uses its address-space, thread and IPC abstractions. Like most OSes, it consists mostly of I/O subsystems (device drivers, protocol stacks, file systems) and resource management. The latter part is particularly small in LionsOS because of its static architecture, which reduces resource management to some simple policy modules that are not only use-case specific (as per our principle) but also local in their nature (e.g. shaping traffic of a network interface).
Some global resource management is required, such as core management (off- and on-lining processor cores based on overall CPU load). This is presently work in progress.
@@ 4.1.1 Device drivers
Applying separation of concerns, we reduce the purpose of a device driver to translate between a hardware-specific device interface and a hardware-independent device class interface: an Ethernet driver does no more than abstracting the specific NIC as a generic Ethernet device.
Unsurprisingly, the Ethernet device-class interface (i.e. the driver’s OS interface) looks similar to that of an actual Ethernet network interface controller (NIC), with some differences that help simplify its use. NICs typically use ring buffers in DMA memory to pass references to DMA buffers from and to the driver; each ring buffer entry contains a pointer to a buffer in the DMA region, together with some meta-data (indicating whether a buffer contains valid data). A NIC usually references two such ring buffers, one for transmit (Tx) and one for receive (Rx) data.
The driver’s OS interface uses buffer queues similar to these hardware-specified ring buffers, but to simplify use we use separate queues for buffers containing valid data and those that do not. This means that the software side of the driver has four queues:
@@
references buffers with valid data provided by the OS for transmission;
@@
These queues are allocated in driver metadata regions, for Ethernet there is one for Tx data (containing the TxA and TxF queues) and a separate one for Rx data (RxA and RxF queues). These are “normal” memory invisible to the device (i.e. not accessed by DMA). For Ethernet we can keep the Tx and Rx regions separate.
Note that the driver only handles pointers to DMA buffers, it has no need to access the actual data. As shown in in the right half of Figure 1, we make this explicit by separating the data region, which contains the buffers to be filled/emptied by the device, from the device metadata region, which contains the HW-defined ring buffers pointing to the data buffers. The data region is not mapped into the driver’s address space, in line with the principle of least privilege. The device control region is mapped to the driver uncached.
In addition there is the device control region, which maps the device registers for memory-mapped I/O. Data and device metadata regions are memory regions accessed by the device via DMA.
@@ 4.1.2 Virtualisers
Drivers abstract the device hardware; in line with separation of concerns they do not deal with sharing the device between multiple clients, and the address translations required for this. This is the responsibility of a separate virtualiser (Virt) component which:
  • shares a physical device between multiple clients;
@@
  • performs cache management (flushing/invalidating) where needed (not required on the x86 architecture, which keeps caches coherent with DMA).
@@
The Virt replicates its driver interface at the client side, meaning each of the Virt’s client-side interfaces looks exactly like its device interface. Specifically it has a metadata region (Tx V MD, Rx V MD) that structurally replicates the respective device metadata region. The key difference is that while the device MD regions use I/O addresses for referring to data buffers, the Virt MD regions refers to data buffers by offsets from the beginning of the respective data region, thus making the Virt’s address-translation task independent of client virtual address-space layout.
The RxVirt needs to inspect the headers of incoming packets so it has the Rx Data region mapped (R/O). The TxVirt does not need to access the data region directly, but may need to have it mapped to perform cache management. For example, the Arm architecture performs cache operations on virtual address ranges, so the TxVirt needs the data region to be mapped into its address space. Arm does not have cache-coherent DMA, so memory that will be transferred to a device using DMA has to be cache-cleaned before DMA occurs. Likewise, the RxVirt on Arm needs to invalidate caches after DMA into its buffers. These mappings are indicated in the left half of Figure 1.
The TxVirt must implement a traffic shaping policy if its clients generate load that exceeds the NIC’s Tx capacity. In line with our principle of use-case specific policy, the Tx policy is as simple as the specific use-case allows, eg. round-robin, priority-based or bandwidth limiting.
The RxVirt at most requires a simple policy: what to do when data arrives for a client whose RxA queue is full. Possible choices are to block, or (more likely) discard the packet and return the buffer to the driver’s RxF queue. We generally avoid this case by ensuring that all client queues are large enough to hold all available Rx buffers, starving the device of buffers if the clients fail to process input fast enough – this leads to the NIC dropping packets under overload without wasting CPU cycles, and leaves the RxVirt policy-free.